OpenStudy (sugarapples):

x 4/5 I need to write in radical form

OpenStudy (tkhunny):

Have you considered the fifth root?

Nnesha (nnesha):

hint:\[\huge\rm \sqrt[n]{x^m} = x^{\frac{ m }{ n } }\]

OpenStudy (sugarapples):



Nnesha (nnesha):

do you know what does they mean by radical form ?? what is radical form ??

OpenStudy (sugarapples):

I don't know

Nnesha (nnesha):

expression that contains radical sign \[\sqrt{}\] square root 4th root etc

Nnesha (nnesha):

radical form should have root so look at the example i gave u before hint:\[\huge\rm \sqrt[n]{x^m} = x^{\frac{ m }{ n } }\]